Monday, Mar. 07: DAY Shift

From Hall A Wiki
Jump to: navigation, search

Go back to Daily Run Plan Daily Run Plans


  • Production data (Kin48_3)

Specific mission

      • from 8AM meeting:
  • Testing the capability of high beam current to Hall-A
  • Put the HOME target/ DVCS HV OFF/ Notify Compton Expert
  • MCC found a several beam trip around RF separator region while increasing higher than 30uA. They estimated several hours work to fix this issue.
  • No success BCM in this time
  • Plan for BCM in this week (after beam study) ?
- The issue of high current is not because of BLM but RF separator which is unusual.
- For now, MCC don't know why beam trip is happened in this region. 
- Therefore, even though they spend couple of hours today for this issue, it would be quite unfruitful.
- Emphasizing the resolving the issue is crucial for our experiment and put the highest priority during the beam study (Thursday).
- PD agreed and will make a plan with Arne.
      • Continue production mode ***

last update Mar 06

  • Polarized beam at 15 uA on LH2 15 cm through Compton chicane !
  • DVCS is running kin_48_3 [here], as off 03-06 5 pm we have accumulate 2.14 C out of our goal of 4.7 C.
  • LHRS is partially detuned: [here]
- LHRS=2.92GeV, Q1/Q2/Q3=1916.640(readout:1900)/1600/1600 (Q1 regul.OFF)
- PreScaleR : PS1=PS2=PS3=1, && PS8=200
- PreScaleL : PS9 =1, + DVCS config.
  • beam locks: no position lock, energy lock in Arc 2: [here]
  • the straw chambers in the LHRS are ON !
  • DVCS daq configuration is [this one], we are measuring about 2% DT at 15uA.
  • Take Compton run every 2 h

    • Automated BTA (aonl4) : Thanks to Bob
- Based on 1/ >2uA, 2/ LHRS and RHRS DAQ status
- please, review ABU and correct if you need before sign off

BPM/Spot++ Check

  • After beam lost for considerable time (like beam study, pass change,...)
  • Ask MCC to run 5uA wth carbon hole target and make sure you can see the hole through Spot++

Please check more info on the white board in the Counting House ~!!